What is Brewster’s angle for light traveling in water that is reflected from glass (n=1.589)?

answer to 3 decimal places


1
Expert's answer
2021-06-01T14:45:07-0400

The special angle of incidence that produces a 90o angle between the reflected and refracted ray is called the Brewster angle, "\u03b8_p."


and "\tan\theta _p=\dfrac{n_2}{n_1}"


here,

"n_1=1\ and\ \\n_2=1.589"


So, "\theta_p\degree=tan^{-1}(\dfrac{1.589}{1})=tan^{-1}(1.589)"


"\theta_p=57.81670025"



Answer:

Brewster's Angle "\boxed{\theta_p= 57.817\degree}"